Chapter 4 Forecasting Homework Assignment

  1. Daily high temperatures in St. Louis for the last week were as follows: 34, 36, 37, 39, 35, 32, and 36 (yesterday).
    1. Forecast the high temperature today, using a 3-day moving average.  
    2. Forecast the high temperature today, using a 2-day moving average.  

For questions c-e use all the days where a 3-day forecast value is possible.

  1. Calculate the mean absolute deviation based on a 3-day moving average.
  2. Compute the mean squared error for the 3-day moving average.
  3. Calculate the mean absolute percent error for the 3-day moving averag

  1. Data collected on the yearly registrations for a Six Sigma seminar at the Quality College are shown in the following table:

Year

1

2

3

4

5

6

Registrations

4500

6500

6775

6225

8075

8520

Forecast (α=0.5)

5000

Forecast (α=0.1)

5000

  1. Use exponential smoothing with a smoothing constant of 0.5 and 0.1 to forecast years 2-6.
  2. What is the MAD value for each forecast (1-6)?
  3. What is the MSE value for each forecast (1-6)?
  4. Which forecast is better? Why?
Answer #1

a)

Forecast of high temperature today, using a 3-day moving average = Average of previous three days high temperature

= (35+32+36)/3

= 34.33

b)

Forecast of high temperature today, using a 2-day moving average = Average of previous three days high temperature

= (32+36)/2

= 34

Parts c) to e) are solved in the following spreadsheet

Mean Absolute Deviation (MAD) = 2.83

Mean Squared Error (MSE) = 10.50

Mean Absolute PErcent Error (MAPE) = 8.17 %

F13 - fx =AVERAGE(F5:F8) @ C A B DE Absolute Squared Deviation Error High temp 3-day MA Absolute Percent Error 1 3.33 2.33 35

EXCEL FORMULAS:

Day High temp 3-day MA Absolute Deviation Squared Error Absolute Percent Error
1 34
2 36
3 37
4 39 =AVERAGE(B2:B4) =ABS(B5-C5) =D5^2 =D5/B5
5 35 =AVERAGE(B3:B5) =ABS(B6-C6) =D6^2 =D6/B6
6 32 =AVERAGE(B4:B6) =ABS(B7-C7) =D7^2 =D7/B7
7 36 =AVERAGE(B5:B7) =ABS(B8-C8) =D8^2 =D8/B8
8 (today) =AVERAGE(B6:B8)
MAD MSE MAPE
=AVERAGE(D5:D8) =AVERAGE(E5:E8) =AVERAGE(F5:F8)